# FedFDP: Federated Learning with Fairness and Differential Privacy

### v1

@@ 7 Privacy Analysis
where $10<\delta<$ 1, $R_{model}^{i}(\alpha)$ is the RDP of $w_{t+1}^{i}$ is computed by Theorem 3 in Appendix A.2, and $R_{loss}^{i}(\alpha$ is the RDP of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ which is computed by Theorem 4 in Appendix A.2.
The details of the proof please refer to Appendix A.2.
@@
Table 1: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$).
@@
Table 2: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$), while the privacy budget $\epsilon=3.52$.
8 Experiments
In this section, we conduct experiments to demonstrate the performance and fairness of FedFair and FedFDP using three real datasets. Additionally, we investigate the impact of various factors on FedFDP, such as fairness hyperparameter $\lambda$, original clipping bound $C$, heterogeneity and scalability. The impact of heterogeneity and scalability please refer to Appendix A.3).
8.1 Default Settings
@@
We conducted experiments on three computer vision (CV) datasets: MNIST LeCun et al. (1998), FashionMNIST Xiao et al. (2017), and CIFAR10 Krizhevsky (2009). And we comparised our algorithms to six baselines, which include FedAvgMcMahan et al. (2017), SCAFFLODKarimireddy et al. (2020), FedProxLi et al. (2020a), FedDynAcar et al. (2021), ALI-DPFLLing et al. (2023) and q-FFLLi et al. (2020b). For the details of parameters settings and the descriptions of baselines please refer to Appendix A.3.1.
8.2 Results
@@
Performance and Fairness. The experimental results of the FedFair algorithm and baseline methods are shown in Table 1. FedFair notably enhances fairness while maintaining substantial model performance across the MNIST, FashionMNIST, and CIFAR10 datasets, with respective increases of 17.0%, 25.7%, and 9.5%. And Table 2 shows the experimental results of the FedFDP algorithm and baseline methods. Under the same privacy budget, FedFDP demonstrates improvements in model accuracy of 1.2%, 1.8%, and 0.4% compared to the best baseline on the three datasets, while also enhancing fairness by 30.3%, 6.7%, and 27.8% respectively compared to the best baseline.
Figure 4: Impact of different $\lambda$ in FedFDF algorithm
@@
Impact of different $\lambda$. We investigated the correlation between $\lambda$ and fairness $\Psi$ in FedFDF algorithm. As depicted in Figure 4, our findings align with the theoretical analysis in Section 6.1: an initial increase in $\lambda$ leads to a decrease in $\Psi$, followed by an increase.
Figure 5: Impact of different $C$ in FedFDF algorithm
@@
Impact of different $C$. In Section 5.1, we emphasized that excessively large or small values of $C$ cannot ensure both performance and fairness simultaneously. As illustrated in Figure 5, we examined the correlation between $C$, performance, and fairness across two datasets, with $C$ taking values of [0.01, 0.1, 1, 5, 10]. Although $\Psi$ reaches its minimum when $C=1$, at this point, due to excessive noise addition, the model does not converge. As $C$ continues to increase, the model’s performance further deteriorates, resulting in an increase in $\Psi$. When $C=0.1$, both accuracy and $\Psi$ achieve relatively good values.
9 Conclusion
@@
In this paper, firstly, we propose a FedFair algorithm effectively addresses fairness concerns by optimizing a novel local loss function. Building upon FedFair, we propose the FedFDP, incorporates a fairness-aware gradient clipping strategy and an adaptive clipping method for additional loss values to achieve both fairness and differential privacy protection. And then, through convergence analysis, we determine the existence of an optimal fairness parameter, $\lambda$, striking a balance between model performance and fairness. We conduct a comprehensive privacy analysis of our approach using RDP and validate our algorithms through extensive experiments. The results indicate that FedFair and FedFDP significantly outperforms state-of-the-art solutions in terms of model performance and fairness. We believe our work contributes a valuable FL framework for addressing fairness and privacy challenges.
@@ Appendix A Appendix
Table 3: Summary of main notations
$F(\mathbf{w}$)
Global loss function
$F_{i}(\mathbf{w}$)
Local loss function for client $i$
$T$
Communication rounds
$t$
Communication round index
$D_{i}$
Dataset of client $i$
$D$
The union of $D_{i}$
$\mathbf{w}_{t}^{i}$
Local model parameter at node $i$ at iteration $t$
$\mathbf{w}^{*}$
Optimal model parameter that minimizes $F(\mathbf{\mathbf{w}})$
$\eta$
Gradient descent step size
$\lambda$
Fairness parameter
$N$
Total number of client
$q$
batch sample ratio
$|\mathcal{B}_{i}|$
Batch size of client $i$, with the mathematical expectation equal to $|D_{i}|\cdot q$
$d$
The dimension of model
$p_{i}$
Weight of client $i$, equals to $|D_{i}|/|D|$
$\xi_{j}$
Single data sample in $D_{i}$, indexed by $j$
$\sigma$
Noise multiplier for gradient
$\sigma_{l}$
Noise multiplier for loss
$C$
Original clipping bound for gradient
$C_{l}$
Original Clipping bound for loss
|||||$\cdot$||| | ⋅ | |
$L_{2}$-norm
A.1 Proof of Theorem 1
@@
$\mathbf{v}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$ (19)
@@
$C^{i,j}_{t}=\min\left(1+\lambda\cdot\Delta^{j}_{loss},\frac{C}{\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}\right).$
In our analysis, we define two virtual sequences $\mathbf{v}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{v}_{t}^{i}$ and $\mathbf{w}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t}^{i}$, which is motivated by Stich [2018]. Therefore,
@@
$\quad-\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$ (20)
A.1.1 Key Lemma
@@ A.1.2 Proof of Lemma 1
Since $\mathbb{E}\mathcal{B}_{0}=0$, we focus on $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$:
$\mathbb{E}\mathcal{B}_{2}\leq\frac{\eta^{2}}{\hat{B}^{2}}\sum_{i=1}^{N}\mathbb{E}\|\sigma C\mathcal{N}(0,\mathbf{I})\|^{2}\leq\frac{\eta^{2}\sigma^{2}C^{2}d}{\hat{B}^{2}},$
@@
$\mathbb{E}\mathcal{B}_{1}\leq\eta^{2}\mathbb{E}\sum_{i=1}^{N}p_{i}\left\|C_{t}^{i}\nabla F_{i}(\mathbf{w}_{t}^{i})\right\|^{2}\leq\eta^{2}C_{t}^{2}G^{2},$
@@
$\mathbb{E}\mathcal{A}_{2}$ $\leq\eta^{2}\left(\frac{\sigma^{2}C^{2}d}{\hat{B}^{2}}+G^{2}C_{t}^{2}\right)$
@@
It’s obvious that $\mathbb{E}\mathcal{C}_{0}=0$.
@@
$\mathbb{E}\mathcal{C}_{1}\leq\sum_{i=1}^{N}p_{i}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\eta^{2}G^{2}C_{t}^{2}$
@@
$\|\nabla F_{i}(\mathbf{w}_{t}^{i})\|^{2}\leq 2L\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}^{*}\right).$ (21)
@@
$\mathbb{E}\mathcal{A}_{1}$ $=\mathbb{E}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\eta^{2}G^{2}C_{t}^{2}+2\eta^{2}C_{t}L\Gamma$
$\quad+C_{t}\mathbb{E}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}-\mu\eta C_{t}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$
$=(1+C_{t})\mathbb{E}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}-\mu\eta C_{t}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$
@@
$\quad\mathbb{E}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}$
$=\mathbb{E}\sum_{i=1}^{N}p_{i}\|(\mathbf{w}_{t}-\mathbf{v}_{t+1}^{i})-(\mathbf{w}_{t}^{i}-\mathbf{v}_{t+1}^{i})\|^{2}$
$\leq\mathbb{E}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{v}_{t+1}^{i}\|^{2}$
@@
$=\mathbb{E}\mathcal{A}_{2}=\eta^{2}C_{t}^{2}G^{2}+\frac{\eta^{2}\sigma^{2}C^{2}d}{\hat{B}^{2}}.$
@@
$\mathbb{E}\mathcal{A}_{1}$ $=(1-\mu\eta C_{t})\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}+(1+C_{t})\eta^{2}C_{t}^{2}G^{2}$
@@
Herterogeneity Scalability
Datasets MNIST FashionMNIST CIFAR10
Method $Dir(0.1)$ $Dir(0.5)$ $Dir(1)$ 10 clients 20 clients 50 clients
Table 4: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability.
Herterogeneity Scalability
Datasets MNIST FashionMNIST CIFAR10
Method $Dir(0.1)$ $Dir(0.5)$ $Dir(1)$ 10 clients 20 clients 50 clients
FedAvg 93.40 |||| 1.1e11 94.67 |||| 4.1e9 75.60 |||| 3.7e9 52.61 |||| 6.1e9 51.58 |||| 2.0e9 55.90 |||| 1.5e9
SCAFFLOD 93.95 |||| 7.0e10 94.06 |||| 1.9e9 72.56 |||| 3.6e9 53.85 |||| 6.2e9 54.34 |||| 2.9e8 53.98 |||| 2.2e9
FedProx 90.50 |||| 5.3e10 94.23 |||| 3.1e9 71.38 |||| 5.6e9 51.34 |||| 4.4e9 53.67 |||| 4.5e8 54.14 |||| 2.2e9
FedDyn 91.42 |||| 5.6e10 93.08 |||| 3.8e9 73.21 |||| 9.8e10 52.14 |||| 5.5e9 50.21 |||| 1.2e10 51.86 |||| 2.0e10
ALI-DPFL 90.89 |||| 3.3e10 93.98 |||| 2.6e9 77.66 |||| 2.5e10 52.05 |||| 3.6e9 38.35 |||| 2.4e10 33.46 |||| 1.6e10
q-FFL 93.74 |||| 7.8e10 93.86 |||| 2.5e8 82.20 |||| 7.4e8 48.46 |||| 4.7e9 51.12 |||| 2.7e8 49.83 |||| 1.1e7
FedFDP 95.13 |||| 2.3e10 94.56 |||| 3.4e9 83.77 |||| 6.0e8 54.21 |||| 2.6e9 54.52 |||| 2.2e8 56.49 |||| 5.3e8
Table 5: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability, while the privacy budget $\epsilon=3.52$.
A.2 Proof of Theorem 2
A.2.1 Privacy Loss of $\mathbf{w}_{t+1}^{i}$
Theorem 3.
$R_{model}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-i}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right),$ (22)
where $\sigma$ is noise multiplier of the $w_{t+1}^{i}$, and $\alpha>1$ is the order.
$proof.$ We will prove this in the following two steps: (i) use the RDP of the sampling Gaussian mechanism to calculate the privacy cost of each model update, and (ii) use the composition of RDP mechanisms to compute the privacy cost of multiple model updates.

### v2

@@ 7.3. Privacy Loss of FedFDP
where $10<\delta<$ 1, $R_{model}^{i}(\alpha)$ is the RDP of $w_{t+1}^{i}$ is computed by Theorem 1, and $R_{loss}^{i}(\alpha)$ is the RDP of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ which is computed by Theorem 2.
Lemma 0.
(Conversion from RDP to DP (Balle et al., 2020)). if a randomized mechanism $f:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$ satisfies $(\alpha,R)$-RDP , then it satisfies$(R+\ln((\alpha-1)/\alpha)-(\ln\delta+\ln\alpha)/(\alpha-1),\delta)$-DP for any $10<\delta<$ 1.
8. Experiments
Table 2. FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$).
@@ 8. Experiments
Table 3. FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$), while the privacy budget $\epsilon=3.52$.
@@
Table 4. FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability.
Herterogeneity Scalability
Datasets MNIST FashionMNIST CIFAR10
Method $Dir(0.1)$ $Dir(0.5)$ $Dir(1)$ 10 clients 20 clients 50 clients
Table 5. FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability, while the privacy budget $\epsilon=3.52$.
Herterogeneity Scalability
Datasets MNIST FashionMNIST CIFAR10
Method $Dir(0.1)$ $Dir(0.5)$ $Dir(1)$ 10 clients 20 clients 50 clients
FedAvg 93.40 |||| 1.1e11 94.17 |||| 4.1e9 85.60 |||| 3.7e9 52.61 |||| 6.1e9 51.58 |||| 2.0e9 55.90 |||| 1.5e9
SCAFFLOD 93.95 |||| 7.0e10 94.06 |||| 1.9e9 84.56 |||| 3.6e9 53.85 |||| 6.2e9 54.34 |||| 2.9e8 53.98 |||| 2.2e9
FedProx 90.50 |||| 5.3e10 94.23 |||| 3.1e9 84.38 |||| 5.6e9 51.34 |||| 4.4e9 53.67 |||| 4.5e8 54.14 |||| 2.2e9
FedDyn 91.42 |||| 5.6e10 93.08 |||| 3.8e9 85.21 |||| 9.8e10 52.14 |||| 5.5e9 50.21 |||| 1.2e10 51.86 |||| 2.0e10
ALI-DPFL 90.89 |||| 3.3e10 93.98 |||| 2.6e9 83.66 |||| 2.5e10 52.05 |||| 3.6e9 38.35 |||| 2.4e10 33.46 |||| 1.6e10
q-FFL 93.74 |||| 7.8e10 93.86 |||| 2.5e8 85.20 |||| 7.4e8 48.46 |||| 4.7e9 51.12 |||| 2.7e8 49.83 |||| 1.1e7
FedFDP 94.13 |||| 2.3e10 94.56 |||| 3.4e9 86.77 |||| 6.0e8 54.21 |||| 2.6e9 54.52 |||| 2.2e8 56.49 |||| 5.3e8
[…]
$\epsilon$ 1.0 2.0 3.0 4.0
Accuracy Fairness Accuracy Fairness Accuracy Fairness Accuracy Fairness
FedAvg 61.68 1.1e6 82.69 2.3e7 84.07 3.8e8 86.83 8.0e8
SCAFFLOD 62.12 1.2e6 82.36 2.3e7 83.39 2.6e9 85.41 5.6e8
FedProx 61.25 1.1e6 81.62 2.3e7 85.32 3.8e8 85.68 2.4e9
FedDyn 62.38 1.5e6 82.43 3.1e7 83.29 2.9e8 84.51 5.2e9
ALI-DPFL 61.78 1.1e6 83.01 2.3e7 84.12 3.9e8 85.58 4.5e9
q-FFL 58.99 3.6e6 80.24 9.6e7 82.92 5.5e8 85.01 3.8e9
FedFDP 63.36 1.0e6 83.15 2.1e7 85.59 9.5e7 86.97 3.8e8
This section outlines our experimental framework, beginning with the presentation of baselines and default parameter configurations. We then detail the experimental assessment of FedFair and FedFDP across dimensions of performance, fairness, heterogeneity, and scalability. Ablation studies focusing on the hyperparameters $\lambda$, $C$ and $\epsilon$ of FedFDP are also included. All tasks will be repeated five times using different random seeds, and the mean and standard deviation will be reported.
8.1. Default Settings
@@
We conducted experiments on three computer vision (CV) datasets: MNIST (LeCun et al., 1998), FashionMNIST (Xiao et al., 2017), and CIFAR10 (Krizhevsky, 2009). And we comparised our algorithms to six baselines:
  1. (1)
    FedAvg(McMahan et al., 2017): FedAvg is the most classic algorithm in federated learning. In fact, we can set $\lambda$ to 0, transforming FedFair and FedFDP into the FedAvg algorithm.
  2. (2)
    SCAFFLOD(Karimireddy et al., 2020): SCAFFOLD uses control variates (variance reduction) to correct for the “client-drift” in its local updates.
  3. (3)
    FedProx(Li et al., 2020b): FedProx introduces a proximal term to the objective function to address the issues of systems heterogeneity and statistical heterogeneity in federated learning.
  4. (4)
    FedDyn(Acar et al., 2021): FedDyn adds linear and quadratic penalty terms to the objective function to solve the problem of inconsistency between the minima of the local-device level empirical loss and those of the global empirical loss.
  5. (5)
    ALI-DPFL(Ling et al., 2023): ALI-DPFL dynamically obtains the optimal number of local DPSGD iterations for clients between any two sequential global updates through minimizing the convergence bound.
  6. (6)
    q-FFL(Li et al., 2020c): q-FFL amplifies loss values for clients with poorer performance by assigning them larger update weights, aiming to improve their model performance and achieve a balanced performance across multiple devices.
For all datasets, we adopted a widely used heterogeneous settings (Lin et al., 2020; Li et al., 2021a) to 10 clients which controlled by a Dirichlet distribution denoted as $Dir(\beta)$, and the default value of $\beta=0.1$ (Lin et al., 2020; Wang et al., 2020). We use a 4-layer CNN architecture (McMahan et al., 2017) which consists of two convolutional layers and two fully connected layers as model architecture.
In hyper-parameters, for FedFair and its baseline algorithms, we set learning rate $\eta=0.1$. For FedFDP and its baseline algorithms, we set the $\eta=1.0$, the batch sample ratio $q=0.05$, clipping bound for gradient $C=0.1$, the noise multipliers for gradient $\sigma=2.0$, the privacy budget $\epsilon=3.52$ and the $\delta=1e-5$. In particular, for the loss values in the FedFDP algorithm that additionally require differential privacy processing, we set the clipping bound for loss $C_{l}=2.5$ and the noise multipliers for loss $\sigma_{l}=5.0$.
8.2. Results
In this subsection, we systematically evaluate the performance and fairness of the FedFair and FedFDP under varying heterogeneity levels, expanding from $Dir(0.1)$ to both $Dir(0.5)$ and $Dir(1)$. We also assess their scalability by scaling the number of clients from the standard 10 to encompass 20 and 50 clients. Furthermore, we examine the influence of three pivotal hyperparameters within the FedFDP algorithm, namely, the adjustments to $\lambda$, $C$ and $\epsilon$.
8.2.1. FedFair
@@
Performance and Fairness. Table 2 presents the experimental results of the FedFair algorithm compared to baseline methods. FedFair significantly enhances fairness while maintaining substantial model performance across the MNIST, FashionMNIST, and CIFAR10 datasets, with respective increases of 17.0%, 25.7%, and 9.6%.
Heterogeneity. Table 4 shows that FedFair outperforms baseline methods in accuracy while maintaining significant fairness when faced with diverse data distributions exhibiting varying levels of heterogeneity. Except for the $Dir(0.5)$ scenario, where it falls slightly short of q-FFL, FedFair demonstrated fairness improvements of 22.0% and 50% over the best baseline in $Dir(0.1)$ and $Dir(1.0)$, respectively.
Scalability. Table 4 indicates that FedFair generally surpasses baselines in accuracy and shows fairness improvements of 9.6%, 15.4%, and 0.01% across three client counts.
8.2.2. FedFDP
Performance and Fairness. Table 3 displays the experimental results of the FedFDP algorithm compared to baseline methods. Under a consistent privacy budget of $\epsilon=3.52$, when accuracy matches the best baseline, our method significantly enhances fairness, with improvements of 30.3%, 6.7%, and 27.8% observed across three separate datasets.
Heterogeneity. Table 5 shows that FedFDP outperforms baseline methods in accuracy while maintaining significant fairness when confronted with diverse data distributions exhibiting varying levels of heterogeneity. Except for the $Dir(0.5)$ scenario, where it slightly underperforms q-FFL, FedFDP shows fairness improvements of 30.3% and 18.9% over the best baseline in $Dir(0.1)$ and $Dir(1.0)$, respectively.
Scalability. Table 5 demonstrates that as the number of clients increases, FedFDP consistently achieves accuracy comparable to the best baseline. Moreover, except in the 50 clients scenario where it slightly lags behind q-FFL in fairness, FedFDP exhibits fairness improvements of 27.8% and 18.5% with 10 and 20 clients, respectively.
Figure 3. Impact of different $\lambda$ in FedFDP algorithm
[…]
Figure 4. Impact of different $C$ in FedFDP algorithm
[…]
@@ 8.2.2. FedFDP
Impact of Different $\lambda$. We investigated the correlation between $\lambda$ and fairness $\Psi$ in the FedFDP algorithm. As depicted in Figure 3, our findings align with the theoretical analysis in Section 6.1: an initial increase in $\lambda$ leads to a decrease in $\Psi$, followed by an increase, indicating the existence of an optimal $\lambda$.
@@
Impact of Different $C$. In Section 5.1, we emphasized that excessively large or small values of $C$ cannot ensure both performance and fairness simultaneously. As illustrated in Figure 4, we examined the correlation between $C$, performance, and fairness across two datasets, with $C$ taking values in the range [0.01, 0.1, 1, 5, 10]. Although $\Psi$ reaches its minimum when $C=1$, the model does not converge due to excessive noise addition. As $C$ continues to increase, the model’s performance further deteriorates, resulting in an increase in $\Psi$. When $C=0.1$, both accuracy and $\Psi$ achieve relatively good performance.
Different $\epsilon$. We conducted experiments with $\epsilon=[1,2,3,4]$ on the FashionMNIST dataset, as detailed in Table 6. For $\epsilon=[1,2,3]$, we allowed larger values of $\epsilon$ to support more iterations. At $\epsilon=4$, we supported a smaller noise multiplier $\sigma=1.65$ with a similar number of iterations as at $\epsilon=3$. The experimental results demonstrate that FedFDP consistently achieves performance on par with the best baseline in terms of accuracy across various privacy budgets. In terms of fairness, the improvements are 9.1%, 8.7%, 67.2%, and 32.1%, respectively.
9. Conclusion
@@
In this paper, a FedFair algorithm is initially proposed to effectively address fairness concerns by optimizing a novel local loss function. The FedFDP, building upon FedFair, is introduced to incorporate a fairness-aware gradient clipping strategy and an adaptive clipping method for additional loss values, thereby achieving both fairness and differential privacy protection. Then, we find an optimal fairness parameter $\lambda^{*}$ through convergence analysis and numerical analysis methods, striking a balance between model performance and fairness. Subsequently, a comprehensive privacy analysis of the approach is conducted using RDP. Through extensive experiments, the results indicate that FedFair and FedFDP significantly outperform state-of-the-art solutions in terms of model performance and fairness. It is believed that our work contributes a valuable FL framework for addressing fairness and privacy challenges.
@@ Appendix A Appendix
A.1. Proof of Theorem 1
@@
(28) $\mathbf{v}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$
@@
$C^{i,j}_{t}=\min\left(1+\lambda\cdot\Delta_{i}^{j},\frac{C}{\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}\right).$
In our analysis, we define two virtual sequences $\mathbf{v}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{v}_{t}^{i}$ and $\mathbf{w}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t}^{i}$, which is motivated by (Stich, 2018). Therefore,
@@
(29) $\quad-\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$
A.1.1. Key Lemma
@@ A.1.2. Proof of Lemma 1
Since $\mathbb{E}[\mathcal{B}_{0}]=0$, we focus on $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$:
$\mathbb{E}[\mathcal{B}_{2}]\leq\frac{\eta^{2}}{\hat{B}^{2}}\sum_{i=1}^{N}\mathbb{E}\|\sigma C\mathcal{N}(0,\mathbf{I})\|^{2}\leq\frac{\eta^{2}\sigma^{2}C^{2}d}{\hat{B}^{2}},$
@@
$\mathbb{E}[\mathcal{B}_{1}]\leq\eta^{2}\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\left\|C_{t}^{i}\nabla F_{i}(\mathbf{w}_{t}^{i})\right\|^{2}\right]\leq\eta^{2}C_{t}^{2}G^{2},$
@@
$\mathbb{E}[\mathcal{A}_{2}]$ $\leq\eta^{2}\left(\frac{\sigma^{2}C^{2}d}{\hat{B}^{2}}+G^{2}C_{t}^{2}\right)$
@@
It’s obvious that $\mathbb{E}[\mathcal{C}_{0}]=0$.
@@
$\mathbb{E}[\mathcal{C}_{1}]\leq\sum_{i=1}^{N}p_{i}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\eta^{2}G^{2}C_{t}^{2}$
@@
(30) $\|\nabla F_{i}(\mathbf{w}_{t}^{i})\|^{2}\leq 2L\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}^{*}\right).$
@@
$\mathbb{E}[\mathcal{A}_{1}]$ $=\mathbb{E}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\eta^{2}G^{2}C_{t}^{2}+2\eta^{2}C_{t}L\Gamma$
$\quad+C_{t}\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}\right]-\mu\eta C_{t}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$
$=(1+C_{t})\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}\right]-\mu\eta C_{t}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$
@@
$\quad\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}\right]$
$=\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|(\mathbf{w}_{t}-\mathbf{v}_{t+1}^{i})-(\mathbf{w}_{t}^{i}-\mathbf{v}_{t+1}^{i})\|^{2}\right]$
$\leq\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{v}_{t+1}^{i}\|^{2}\right]$
@@
$=\mathbb{E}\left[\mathcal{A}_{2}\right]=\eta^{2}C_{t}^{2}G^{2}+\frac{\eta^{2}\sigma^{2}C^{2}d}{\hat{B}^{2}}.$
@@
$\mathbb{E}\left[\mathcal{A}_{1}\right]$ $=(1-\mu\eta C_{t})\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}+(1+C_{t})\eta^{2}C_{t}^{2}G^{2}$
@@
A.2. Proof of Theorem 1
$proof.$ We will prove Theorem 1 in the following two steps: (i) use the RDP of the sampling Gaussian mechanism to calculate the privacy cost of each model update, and (ii) use the composition of RDP mechanisms to compute the privacy cost of multiple model updates.